The business customer Annual Net Income was $85,000,000. the loan approval says that the loan will have a 5-year maturity, based upon a 25-year amortization, at a rate of 5.25%, with a 25% down payment, and the DSC must be 1.25x the Quarterly Payments. What is the maximum building cost that the customer can afford under the same terms as above?
Answer #1

Quarterly Income = 85,000,000/4 = $21,250,000

Maximum Quarterly Payment = 21,250,000/1.25 = $17,000,000

Quarterly Installment = P×[{R×(1+R)n}/{(1+R)n -1}]

where, P=Principal Amount

R=Interest Rate/4 = 5.25/4 = 1.3125%

  N=number of payments = 25 × 4 = 100

Therefore, 17,000,000 =P×[{0.01325×(1+0.01325)100}/{(1+0.01325)100-1}]

17,000,000=P×(0.04941/2.7296)

Therefore, P = $939,145,921.88

Maximum Loan Amount = $939,145,921.88

Therefore, Maximum Building Cost will be 939,145,921.88/75% = $1,252,194,562.5